### v1

# On vanishing of higher direct images of the structure sheaf

Abstract.
We show the vanishing of the first direct image of the structure sheaf of a normal excellent scheme $X$ which is mapped properly and birationally over an excellent regular scheme of any dimension. On the other hand, for any dimension greater than two, we show examples of a proper birational morphism from a normal and Cohen-Macaulay scheme to a regular scheme such that the second direct image does not vanish and has an isolated support.
1. Introduction
Let $\varphi:X\to Y:={\mathrm{Spec\ }}R$ be a projective birational morphism of schemes. It is well known that if $X$ and $Y$ are regular schemes of finite type over a field of characteristic 00 the following holds:
$R^{i}\varphi_{*}{\mathcal{O}}_{X}=0\ \mbox{ for\ every}\ i>0$
by [7, pp.144-145]. This is generalized by [3, Theorem 1.1] for excellent regular schemes $X,Y$.
To illustrate a history, we temporarily assume that $Y$ is a variety over a field $k$ of characteristic 00 with singularities. Let $\varphi:X\to Y$ be a resolution of the singularities of $Y$. Then $R^{i}\varphi_{*}{\mathcal{O}}_{X}$ ($i\in{\mathbb{N}}$) play important roles in measuring how far the singularities of $Y$ from regularity since these are zero if $Y$ is regular. The closest singularities to a regular point in this viewpoint are rational singularities defined as follows:
Definition 1.1.
Assuming ${\operatorname{char}}k=0$, we say that $Y$ has rational singularities if for every resolution $\varphi:X\to Y$ of the singularities of $Y$, the following holds:
(1) $R^{i}\varphi_{*}{\mathcal{O}}_{X}=0\ \mbox{for\ every\ }i>0.$
Here, note that once the vanishings above hold for a resolution $\varphi$, then the vanishings also hold for every resolution of the singularities of $Y$. It is also notable that a rational singularity automatically has the Cohen-Macaulay (CM for short) property, which is considered an essential condition for a singularity to be “good”. On the other hand, in positive characteristic, the vanishing (1) does not automatically imply CM property. Based on this and the present situation, the existence of resolutions of singularities in positive characteristic is not yet proved, there is a challenge reasonably defining a rational singularity without using a resolution of the singularities as follows:
Definition 1.2 ([11, Definition 1.3]).
A scheme $Y$ is said to have rational singularities, if
  1. (i)
    Y is an excellent normal Cohen-Macaulay scheme that admits a dualizing complex, and
  2. (ii)
    for each excellent normal Cohen-Macaulay scheme $X$, and each $\varphi:X\to Y$ locally projective birational morphism, the following vanishing holds:
    $R^{i}\varphi_{*}{\mathcal{O}}_{X}=0\ \mbox{for\ every\ }i>0$
From now on, we distinguish rational singularities in ${\rm char}k=0$ and that in the definition 1.2 by denoting rational singularities and “rational singularities”, respectively. Definition 1.2 asserts that $\varphi:X\to Y$ plays the role of a resolution of the singularities to characterize “rational singularities”. If the definition of “rational singularity” is reasonable, in principle, a regular point should be “rational” (cf. [11, Theorem 9.12]). However, Linquan Ma shows an example of a regular point that does not satisfy this definition of “rational singularity”. More concretely, there is a projective birational morphism $\varphi:X\to{\mathbb{A}}^{3}_{\mathbb{C}},$ from normal and CM variety $X$, such that $R^{2}\varphi_{*}{\mathcal{O}}_{X}\neq 0$ (see Example 3.2). It means that $\varphi:X\to Y$ in Definition 1.2 does not play the role of a resolution of the singularities of $Y$ to characterize rational singularities.
Here, note that in this example, we have $N=3$, then $N-1=2$. So, for $N\geq 4$ one may still expect that $\varphi:X\to Y$ plays the role of a resolution of the singularities to characterize an isolated CM singularity. Remember that, in ${\rm char}k=0$, if $\varphi:X\to Y$ is a resolution of isolated CM singularity (therefore $R^{i}\varphi_{*}{\mathcal{O}}_{X}|_{Y\setminus\{0\}}=0$ $(i>0)$ holds for the singular point $0\in Y$) of $N$-dimensional $Y$, then, $\ R^{i}\varphi_{*}{\mathcal{O}}_{X}=0\ \mbox{ holds\ for\ every\ }0<i<N-1.$
Expectation. For a CM closed point $0\in Y$ on an $N$-dimensional scheme ($N\geq 3$) and a proper birational morphism $\varphi:X\to Y$ from a normal CM scheme $X$, if $R^{i}\varphi_{*}{\mathcal{O}}_{X}|_{Y\setminus\{0\}}=0$ $(i>0)$ holds, then the following vanishings holds,
(2) $\ R^{i}\varphi_{*}{\mathcal{O}}_{X}=0\ \mbox{ holds\ for\ every\ }0<i<N-1.$
In this paper, we show that this expectation is satisfied only partially. Concretely, if $i=1$, the expectation is satisfied in a stronger form (Theorem 1.3), and as its application we obtain the vanishing
$\ R^{1}\varphi_{*}{\mathcal{O}}_{X}=0$
for regular $Y$ (Corollary 1.4). Here, we do not need CM for $X$. If $i=2$, the expectation does not hold even for smooth $Y$ and normal CM variety $X$ of any dimension $N\geq 3$. We summarize the results in the following.
Theorem 1.3.
Let $Y$ be an excellent scheme with the Serre’s conditions $R_{2}$ and $S_{3}$ of dimension $N\geq 1$. Let $\varphi:X\to Y$ be a proper birational morphism such that $X$ is excellent and normal. Then, the following vanishing holds,
(3) $\ R^{1}\varphi_{*}{\mathcal{O}}_{X}=0.$
The following corollary follows immediately from Theorem 1.3.
Corollary 1.4.
Let $Y$ be an excellent regular scheme of dimension $N\geq 1$. Let $\varphi:X\to Y:={\mathrm{Spec\ }}R$ be a proper birational morphism with excellent and normal $X$. Then
$R^{1}\varphi_{*}{\mathcal{O}}_{X}=0.$
Here, note that we assume only normality on $X$.
Example 1.5.
Let $(R,{\mathfrak{m}})$ be a regular local ring of dimension $N\geq 3$ essentially of finite type over ${\mathbb{C}}$. There exists a projective birational morphism $\varphi:X\to Y:={\mathrm{Spec\ }}R$ with a normal and CM variety $X$ such that
$R^{2}\varphi_{*}{\mathcal{O}}_{X}\neq 0,\ \ \mbox{and}\ \ R^{i}\varphi_{*}{\mathcal{O}}_{X}|_{Y\setminus\{{\mathfrak{m}}\}}=0\ \ (i>0)$
This paper is organized as follows: in the second section we prove the theorem and the third section we prove the example.
Acknowledgement The authors express their heartfelt thanks to Sándor Kovács, Linquan Ma, Mircea Mustaţǎ, Karl Schwede, Kohsuke Shibata, Shunsuke Takagi for useful informations and discussions, in particular Linquan Ma’s kindness to allow us including his example (Example 3.2) in this paper.
2. Vanishing of $R^{1}$
In this section, we concentrate to prove Theorem 1.3. We use the same symbols as in Theorem 1.3. Let $\varphi:X\to Y$ be as in the theorem. As our problem is local, we may assume that $Y={\mathrm{Spec\ }}R$, where $(R,{\mathfrak{m}})$ is an excellent local ring satisfying $R_{2}$ and $S_{3}$. As $Y$ is affine, we sometimes write $H^{i}(X,{\mathcal{O}}_{X})$ for $R^{i}\varphi_{*}{\mathcal{O}}_{X}$. The following lemma will allow us to assume that $\varphi$ is a projective birational morphism.
Lemma 2.1.
Let $\varphi:X\to Y$ be a proper birational morphism with normal $X$. Assume there exists a proper birational morphism $\psi:X^{\prime}\to X$ such that $R^{1}(\varphi\circ\psi)_{*}{\mathcal{O}}_{X^{\prime}}=0$. Then,
$R^{1}\varphi_{*}{\mathcal{O}}_{X}=0.$
Proof.
Consider the Leray spectral sequence
$E_{2}^{p,q}=R^{p}\varphi_{*}(R^{q}\psi_{*}{\mathcal{O}}_{X^{\prime}})\Rightarrow E^{p+q}=R^{p+q}(\varphi\circ\psi)_{*}{\mathcal{O}}_{X^{\prime}}$
induces the exact sequence
$0\to R^{1}\varphi_{*}(\psi_{*}{\mathcal{O}}_{X^{\prime}})\to R^{1}(\varphi\circ\psi)_{*}{\mathcal{O}}_{X^{\prime}}\to\varphi_{*}(R^{1}\psi_{*}{\mathcal{O}}_{X^{\prime}})\to.$
As the middle term is zero, the first term becomes zero too, which implies $R^{1}\varphi_{*}{\mathcal{O}}_{X}=0$ by the normality of $X$. ∎
In our situation of Theorem 1.3, apply Chow’s lemma ([6, II, Theorem 5.6.1]) to the given proper birational morphism $\varphi:X\to Y$, we obtain a birational morphism $\psi:X^{\prime}\to X$ such that the composite $\Phi=\varphi\circ\psi:X^{\prime}\to Y$ is projective birational. By Lemma 2.1, for the proof of $R^{1}\varphi_{*}{\mathcal{O}}_{X}=0$ it is sufficient to show $R^{1}\Phi_{*}{\mathcal{O}}_{X^{\prime}}=0$. Now we may assume that $\varphi:X\to Y={\mathrm{Spec\ }}R$ is projective birational and $X$ is normal. Then, there is an ideal $I\subset R$ whose blow-up gives $\varphi$. Then, we can see that $00ptI>$ 0. Denote the Rees algebra of $I$ by
${\mathcal{R}}(I):=R\oplus IT\oplus I^{2}T^{2}\oplus\cdots,$
where $T$ is an indeterminate, and denote the fiber algebra by
$G(I):=R/I\oplus(I/I^{2})T\oplus(I^{2}/I^{3})T^{2}\oplus\cdots.$
We denote the maximal ideal of ${\mathcal{R}}(I)$ by
${\mathcal{M}}:={\mathfrak{m}}\oplus IT\oplus I^{2}T^{2}\oplus\cdots.$
Note that
$X={\mathrm{Proj\ }}{\mathcal{R}}(I)\simeq{\mathrm{Proj\ }}{\mathcal{R}}(I^{n}),$
for every $n\in{\mathbb{N}}$. An ideal $I\subset R$ is called a normal ideal if every $I^{n}$ $(n\in{\mathbb{N}})$ is integrally closed. We know that $I\subset R$ is normal if and only if ${\mathcal{R}}(I)$ is a normal ring.
We will interpret our vanishing problem into the depth of the Rees algebra in the following lemma, where a local ring $(R,{\mathfrak{m}})$ is called equidimensional if $\dim R/\mathfrak{p}=\dim R$ holds for every minimal prime ideal $\mathfrak{p}$ of $R$.
Lemma 2.2.
Let $(R,{\mathfrak{m}})$ be an excellent equidimensional local ring of $0ptR\geq 3$. Let $I\subset R$ be an ideal with ${\mathrm{grade}}I>0$. Suppose that $X={\mathrm{Proj\ }}\mathcal{R}(I)$ is normal. Then, for $n\gg 0$ the following holds::::
$0pt_{\mathcal{M}}\mathcal{R}(I^{n})\geq 3$
i.e., $H_{\mathcal{M}}^{i}(\mathcal{R}(I^{n}))=0$, for $i\leq 2$.
Proof.
Since $X={\mathrm{Proj\ }}\mathcal{R}(I)$ is normal, all large power $I^{n}$ is normal by [10, Proposition 3.6]. So we may assume that $I$ is normal. This implies that $0pt_{{\mathcal{M}}}G(I^{n})\geq 2$ for large $n\gg 0$ from [10, Theorem 3.4] (cf. [8, Theorem 3.1]).
Fix an integer $n$ such that $0pt_{{\mathcal{M}}}G(I^{n})\geq 2$. If $0pt_{\mathcal{M}}G(I^{n})\geq 3$, then
$0pt_{{\mathcal{M}}}\mathcal{R}(I^{n})\geq 0pt_{{\mathcal{M}}}G(I^{n})\geq 3$
by [9, Proposition 3.6]. Otherwise, $0pt_{{\mathcal{M}}}G(I^{n}){=2<0ptR}$. It follows from [9, Theorem 3.10] that
$30pt_{{\mathcal{M}}}\mathcal{R}(I^{n})=0pt_{{\mathcal{M}}}G(I^{n})+1=2+1=$ 3 .
If $N=\dim Y=1$, then the vanishing $R^{1}\varphi_{*}{\mathcal{O}}_{X}=0$ is trivial.
If $N=2$, $R_{2}$ implies that $Y$ is regular, and by Lipman’s theorem ([12]), there exists a resolution $\psi:X^{\prime}\to X$ of the singularities of $X$. Then, applying [3, Theorem 1.1] to the morphism $\varphi\circ\psi:X^{\prime}\to Y$, we obtain $R^{1}(\varphi\circ\psi)_{*}{\mathcal{O}}_{X^{\prime}}=0$ which yields
$R^{1}\varphi_{*}{\mathcal{O}}_{X}=0$
by Lemma 2.1.
For $N\geq 3$, note that we can apply Lemma 2.2 to our case in Theorem 1.3 since $Y={\mathrm{Spec\ }}R$ is the surjective image of normal scheme $X={\mathrm{Proj\ }}{\mathcal{R}}(I)$. Suppose that $R^{1}\varphi_{*}{\mathcal{O}}_{X}\neq 0$ and let $C\subset Y$ be an irreducible component of ${\mathrm{supp}}R^{1}\varphi_{*}{\mathcal{O}}_{X}$. Let ${\mathfrak{p}}\subset R$ be the ideal defining $C$, then, $\dim R_{\mathfrak{p}}\geq 3$ by the discussions for $N\leq 2$. We also have $0ptR_{\mathfrak{p}}\geq 3$ as we assume $S_{3}$ for $R$. In this setting we will prove:
$H^{1}(\varphi^{-1}({\mathrm{Spec\ }}R_{\mathfrak{p}}),{\mathcal{O}}_{{\mathrm{Spec\ }}R_{\mathfrak{p}}})=0,$
which is a contradiction to the definition of ${\mathfrak{p}}$, and therefore, completes the proof of the theorem. Denote $R_{\mathfrak{p}}$ by $R$ and denote its maximal ideal by ${\mathfrak{m}}$. Replace $Y$ and $X$ accordingly. Let $E:=\varphi^{-1}({\mathfrak{m}})$, and
$\varphi^{\prime}:X\setminus E\to Y\setminus\{{\mathfrak{m}}\}$
the restriction of $\varphi:X\to Y$. Then, by the construction of $R$, we have the following vanishing
(4) $R^{1}\varphi^{\prime}_{*}{\mathcal{O}}_{X\setminus E}=0.$
For the proof of the theorem, we show the following isomorphism first:
(5) $H^{1}(X,{\mathcal{O}}_{X})\simeq H^{1}_{E}(X,{\mathcal{O}}_{X}).$
For (5), consider the exact sequence:
$H^{0}(X,{\mathcal{O}}_{X})\stackrel{{\scriptstyle\alpha}}{{\longrightarrow}}H^{0}(X\setminus E,{\mathcal{O}}_{X})\to H^{1}_{E}(X,{\mathcal{O}}_{X})\to H^{1}(X,{\mathcal{O}}_{X})\to H^{1}(X\setminus E,{\mathcal{O}}_{X}).$
For the required isomorphism, it is sufficient to prove the following (I) and (II):
(I) $H^{1}(X\setminus E,{\mathcal{O}}_{X})=0$, and (II) $\alpha$ is surjective.
Indeed, for the morphism $\varphi^{\prime}:X\setminus E\to Y\setminus\{{\mathfrak{m}}\}$ we have the spectral sequence
$E_{2}^{p,q}=H^{p}(Y\setminus\{{\mathfrak{m}}\},R^{q}\varphi^{\prime}_{*}({\mathcal{O}}_{X\setminus E}))\Rightarrow E^{p+q}=H^{p+q}(X\setminus E,{\mathcal{O}}_{X}).$
In the canonical exact sequence
$0\to E_{2}^{1,0}\to E^{1}\to E_{2}^{0,1},$
$E_{2}^{1,0}=H^{1}(Y\setminus\{{\mathfrak{m}}\},{\mathcal{O}}_{Y})=H_{\mathfrak{m}}^{2}(R)=0$ since $0ptR\geq 3$, while $E_{2}^{0,1}=H^{0}(Y\setminus\{{\mathfrak{m}}\},R^{1}\varphi^{\prime}_{*}{\mathcal{O}}_{X\setminus E})=0$ by (4). This gives the vanishing $E^{1}=H^{1}(X\setminus E,{\mathcal{O}}_{X})=0$ as claimed in (I).
On the other hand, the above spectral sequence satisfies $E_{2}^{p,q}=0$ for $p<0$ or $q<0$, which gives the isomorphism $E^{0}=E_{2}^{0,0}$. Then, the homomorphism
$\alpha:H^{0}(X,{\mathcal{O}}_{X})\to H^{0}(X\setminus E,{\mathcal{O}}_{X})=E^{0}$
is the same as
$H^{0}(Y,{\mathcal{O}}_{Y})\to H^{0}(Y\setminus\{{\mathfrak{m}}\},{\mathcal{O}}_{Y})=E_{2}^{0,0}.$
Hence, $\alpha$ is isomorphic because of $H_{{\mathfrak{m}}}^{i}(R)=0$ for $i=0,1$. Q.E.D. of (II).
Next, consider the (SdS)-sequence
$\to H_{{\mathfrak{m}}}^{1}(R)\to H_{E}^{1}(X,{\mathcal{O}}_{X})\to H_{\mathcal{M}}^{2}({\mathcal{R}}(I^{n}))_{0}\to$
where $H_{\mathcal{M}}^{2}({\mathcal{R}}(I^{n}))_{0}$ is the degree 0 part of the graded module $H_{\mathcal{M}}^{2}({\mathcal{R}}(I^{n}))$. Here, $H_{{\mathfrak{m}}}^{1}(R)=0$ follows from $0ptR\geq 3$. On the other hand, $H_{\mathcal{M}}^{2}({\mathcal{R}}(I^{n}))_{0}=0$ for $n\gg 0$ follows from Lemma 2.2. Then, we obtain $H_{E}^{1}(X,{\mathcal{O}}_{X})=0$ which yields $H^{1}(X,{\mathcal{O}}_{X})=0$ by (5). [QED of Theorem 1.3]
The following is a well-known example that $H^{1}(X,{\mathcal{O}}_{X})={\mathbb{C}}$ for the minimal resolution $\varphi:X\to Y={\mathrm{Spec\ }}R$ for which the conditions of Theorem 1.3 do not hold. We can show that the ideal which gives $\varphi$ is normal. We also have the depth of the Veronese rings of the Rees algebra.
Example 2.3.
Let $R=\mathbb{C}[[x,y,z]]/(x^{3}+y^{3}+z^{3})$. Then, the following hold:
  1. (1)
    $R$ is a normal domain with $20ptR=\dim R=$ 2.
  2. (2)
    ${\mathfrak{m}}=(x,y,z)R$ is normal.
  3. (3)
    $\big{[}H^{2}_{\mathcal{M}}(\mathcal{R}({\mathfrak{m}}^{n}))\big{]}_{0}\neq 0$ for every integer $n\geq 1$. In particular,
    $0pt_{\mathcal{M}}\mathcal{R}({\mathfrak{m}}^{n})=2\ \ \mbox{for\ every\ }n\geq 1$
3. Non-vanishing of $R^{2}$
An example of non-vanishing $R^{2}f_{*}{\mathcal{O}}_{X}\neq 0$ for 3-dimensional case was known in the last century for normal and non-CM $X$.
Example 3.1 (S.D. Cutkosky [4, Section 3]).
There exists a projective birational morphism $f:Z\to{\mathbb{A}}_{\mathbb{C}}^{3}$ with normal $Z$, isomorphic outside the origin $0\in{\mathbb{A}}_{\mathbb{C}}^{3}$ such that
$R^{2}f_{*}{\mathcal{O}}_{Z}\neq 0.$
It is constructed as follows:
First, blowup $X_{1}\to{\mathbb{A}}_{\mathbb{C}}^{3}$ at the origin $\{0\}$, then take an elliptic curve $C$ in the exceptional divisor $E_{1}\simeq{\mathbb{P}}_{\mathbb{C}}^{2}$. Then, blowup $X_{2}\to X_{1}$ at appropriate 12 points in $C$. Denote the proper transform of $C$ in $X_{2}$ by $C^{\prime}$. Then, there exists a base point free linear system on $X_{2}$ which gives a projective birational morphism $\psi:X_{2}\to Z$ which contracts $C^{\prime}$ to a normal point in $Z$. The point $P=\psi(C^{\prime})$ is not a CM point of $Z$. Let $f:Z\to{\mathbb{A}}_{\mathbb{C}}^{3}$ be the canonical morphism. Then, by the Leray’s spectral sequence, we obtain $R^{2}f_{*}{\mathcal{O}}_{Z}\neq 0.$
The following is an example that a projective birational morphism $\varphi:X\to{\mathbb{A}}_{\mathbb{C}}^{3}$ such that $X$ is normal and CM but $R^{2}\varphi_{*}{\mathcal{O}}_{X}\neq 0$.
Example 3.2 (L. Ma).
Let $f:Z\to{\mathbb{A}}_{\mathbb{C}}^{3}$ be Cutkosky’s example 3.1. Let
$g:X\to Z$
be a projective birational morphism such that $X$ is arithmetically normal and arithmetically Cohen-Macaulay as in [1, Corollary 1.4] (in particular, $X$ is normal and Cohen-Macaulay). We first claim that
$Rg_{*}{\mathcal{O}}_{X}={\mathcal{O}}_{Z}$
[…]
Now consider the projective birational morphism $\varphi:=f\circ g:X\to{\mathbb{A}}_{\mathbb{C}}^{3}$. We have $X$ is normal and Cohen-Macaulay, and $R(f\circ g)_{*}{\mathcal{O}}_{X}=Rf_{*}Rg_{*}{\mathcal{O}}_{X}=Rf_{*}{\mathcal{O}}_{Z}\neq{\mathcal{O}}_{{\mathbb{A}}_{\mathbb{C}}^{3}}$. Concretely, $R^{2}(f\circ g)_{*}{\mathcal{O}}_{X}\neq 0$.
Before stepping into the construction of our example, we explain briefly about Brodmann’s Macaulayfications [1, Corollary 1.4] which is used in Example 3.2, and also about [2, Corollary 6.12] which will be used in our example.
The Corollary 1.4 in[1] studies a three-dimensional quasi-projective variety $V$ over an algebraically closed field. Under the assumption that $V$ has only finite number of isolated non-CM points, the corollary states that there exists a reduced (possibly non-irreducible) curve passing through these points such that the blowup at the curve gives arithmetically normal and arithmetically CMfication of $V$. This process preserves normality and regularity, i.e., any point of the fiber of a normal (resp. regular) point is also normal (resp. regular).
The Corollary 6.12 in [2] study an arbitrary dimensional local ring $R$. Under the assumption that ${\mathrm{Spec\ }}R$ is CM on ${\mathrm{Spec\ }}R\setminus\{{\mathfrak{m}}\}$, the Corollary 6.12 states that there exists a curve passing through the point ${\mathfrak{m}}$ such that the blowup at the curve gives arithmetically normal and arithmetically CMfication of ${\mathrm{Spec\ }}R$. This process also preserves normality and regularity.
[Construction and the proof of Example 1.5] We use Example 3.2 and denote $Z$ and $X$ in Example 3.2 by $Z^{(3)}$ and $X^{(3)}$, respectively to clarify their dimensions. Let $f^{(3)}:Z^{(3)}\to{\mathbb{A}}_{\mathbb{C}}^{3}$ be the Cutkosky’s morphism. We also denote the blowup $X^{(3)}\to Z^{(3)}$ at $C\subset Z^{(3)}$ by $g^{(3)}$ and the composite $f^{(3)}\circ g^{(3)}:X^{(3)}\to{\mathbb{A}}_{\mathbb{C}}^{3}$ by $\varphi^{(3)}$. Then, $\varphi^{(3)}$ is isomorphic outside 00 if and only if $C\subset{f^{(3)}}^{-1}(0)$, and otherwise $\varphi^{(3)}$ is isomorphic outside a curve in ${\mathbb{A}}_{\mathbb{C}}^{3}$. As $X^{(3)}$ is projective birational over ${\mathbb{A}}_{\mathbb{C}}^{3}$, there exists an ideal $I\subset{\mathbb{C}}[x_{1},x_{2},x_{3}]$ such that
$X^{(3)}={\mathrm{Proj\ }}{\mathcal{R}}(I).$
Corollary 1.4 in [1] guarantees the existence of a curve $C\subset Z^{(3)}$ whose blowup gives $X^{(3)}$. However, it is not clear if the curve is in the fiber ${f^{(3)}}^{-1}(0)$ of the origin $0\in{\mathbb{A}}_{\mathbb{C}}^{3}$ or not.
So we note that the following two cases can happen, where we note that $f^{(3)}$ is isomorphic outside the origin $0\in{\mathbb{A}}_{\mathbb{C}}^{3}$ and
  1. (a)
    The ideal $I$ is a primary ideal associated to $(x_{1},x_{2},x_{3})$, in case $C\subset{f^{(3)}}^{-1}(0)$, and
  2. (b)
    The ideal $I$ has height $2$ associated primes which define a smooth (possibly non-irreducible) curve on ${\mathbb{A}}_{\mathbb{C}}^{3}\setminus\{0\}$, in case $C\not\subset{f^{(3)}}^{-1}(0)$.
Here, about (a), we note that an ideal $I$ whose blowup gives an isomorphism outside a point is locally principal there in general. However, as we are working on UFD ${\mathbb{C}}[x_{1},x_{2},x_{3}]$ we can replace $I$ by a maximal primary ideal.
Starting with the three-dimensional model $\varphi^{(3)}:X^{(3)}={\mathrm{Proj\ }}R(I)\to{\mathbb{A}}_{\mathbb{C}}^{3}$, we construct $N$-dimensional model $\varphi^{(N)}:X^{(N)}\to{\mathbb{A}}_{\mathbb{C}}^{N}$ by induction on the dimension $N$. Then, we will show that $X^{(N)}$ is normal and CM, and $H^{2}(X^{(N)},{\mathcal{O}}_{X^{(N)}})\neq 0$ and also $R^{i}\varphi^{(N)}{\mathcal{O}}_{X^{(N)}}|_{{\mathbb{A}}_{\mathbb{C}}^{N}\setminus\{0\}}=0\ \ (i>0)$.
Lemma 3.3.
Let $S_{N}:={\mathbb{C}}[x_{1},\ldots,x_{N}]$ be the polynomial ring of $N$-variables. In the connection between $S_{N}$ and $S_{N+1}$, we denote the $N+1$-th variable $x_{N+1}$ by $z$. Then, $S_{N+1}=S_{N}[z]$ and $S_{N}$ is a direct summand of $S_{N+1}$. Let $I\subset S_{N}$ be a non-zero ideal and consider the Rees algebra of $(I,z)\subset S_{N+1}$ over $S_{N+1}$. Then, we obtain the following:
  1. (i)
    ${\mathcal{R}}(I)$ is a direct summand of ${\mathcal{R}}((I,z))$.
  2. (ii)
    If ${\mathcal{R}}(I)$ is normal, then ${\mathcal{R}}((I,z))$ is also normal.
  3. (iii)
    If $X^{(N)}:={\mathrm{Proj\ }}{\mathcal{R}}(I)$ is normal, then for $n\gg 0$, $Z^{(N+1)}:={\mathrm{Proj\ }}{\mathcal{R}}((I^{n},z))$ is normal.
Proof.
Every element of ${\mathcal{R}}((I,z))$ is uniquely expressed as the sum of an element without $z$ and that with $z$. The element without $z$ is contained in ${\mathcal{R}}(I)$, this yields the proof of (i).
For the proof of (ii), it is enough to prove that $(I,z)^{n}$ is integrally closed by an induction on $n\geq 1$.
Now suppose $f(z)=f_{0}+f_{1}z+\cdots+f_{\ell}z^{\ell}\in\overline{(I,z)^{n}}$, where $f_{i}\in R$. We first show $f_{0}\in\overline{I^{n}}=I^{n}$. By definition, there exist $N\geq 1$ and $a_{i}(z)\in(I,z)^{ni}$ for $i=1,2,\ldots,N$ such that
(6) $f(z)^{N}+a_{1}(z)f(z)^{N-1}+\cdots+a_{N}(z)=0,$
where $a_{i}(z)=a_{i0}+a_{i1}z+\cdots+a_{i\ell_{i}}z^{\ell_{i}}\in(I,z)^{ni}$ for each $i$. In the degree 00 term in Eq(6), we have
$f_{0}^{N}+a_{10}f_{0}^{N-1}+\cdots+a_{N0}=0$
For each $i$, $a_{i}(z)\in(I,z)^{ni}$ implies $a_{i0}\in I^{ni}$. Hence the above equation gives an integral equation of $f_{0}\in\overline{I^{n}}=I^{n}$, as required.
In particular, when $n=1$, this means $f=f_{0}+z(f_{1}+f_{2}z+\cdots+f_{\ell}z^{\ell-1})\in(I,z)$.
Suppose $n\geq 2$ and $\overline{(I,z)^{k}}=(I,z)^{k}$ for every $k=1,2,\ldots,n-1$. Then we must show $\overline{(I,z)^{n}}=(I,z)^{n}$. Now suppose $f(z)\in\overline{(I,z)^{n}}$. By argument as above, we obtain $f_{0}\in I^{n}$. So we may assume $f_{0}=0$. Then we can write $f(z)=z\cdot g(z)$. Notice that $g(z)\in\overline{(I,z)^{n}}\colon z=\overline{(I,z)^{n-1}}$. By assumption, we get $f(z)=z\cdot g(z)\in z(I,z)^{n-1}\subset(I,z)^{n}$.
(iii) By the assumption of (iii) and also by $00ptI>$ 0, we see that $R(I^{n})$ is normal for $n\gg 0$ ([10, Proposition 3.6]) Then, by (ii) $R((I^{n},z))$ is also normal for $n\gg 0$, a fortiori $Z^{(N+1)}:={\mathrm{Proj\ }}{\mathcal{R}}((I^{n},z))$ is normal. ∎
Once we have $\varphi^{(N)}:X^{(N)}={\mathrm{Proj\ }}{\mathcal{R}}(I)\to{\mathbb{A}}_{\mathbb{C}}^{N}$ for $I\subset S_{N}$ such that $X^{(N)}$ is normal and CM with $H^{2}(X^{(N)},{\mathcal{O}}_{X^{(N)}})\neq 0$ then we obtain the $(N+1)$-dimensional model
$f^{(N+1)}:Z^{(N+1)}={\mathrm{Proj\ }}{\mathcal{R}}((I^{n},z))\to{\mathbb{A}}_{\mathbb{C}}^{N+1}$
for $n\gg 0$ such that
$H^{2}(Z^{(N+1)},{\mathcal{O}}_{Z^{(N+1)}})\neq 0,$
by (i) in Lemma 3.3. As $X^{(N)}$ is normal, so is $Z^{(N+1)}$ by (iii) of Lemma 3.3. However even if $X^{(N)}$ is CM, the scheme $Z^{(N+1)}$ is not necessarily CM. We will see that $Z^{(N+1)}$ has at most one non-CM point so that we can apply Brodmann’s Macaulayfication to get $X^{(N+1)}$ which will be normal and CM satisfying
$H^{2}(X^{(N+1)},{\mathcal{O}}_{X^{(N+1)}})\neq 0.$
Lemma 3.4.
Let $I\subset S_{N}:={\mathbb{C}}[x_{1},\ldots,x_{N}]$ be an ideal either of the following types:
  1. (a)
    a maximal primary associated to $(x_{1},\ldots,x_{N})\subset S_{N}$, or
  2. (b)
    $\surd\overline{I}$ defines a curve which is smooth (possibly non-irreducible) on ${\mathbb{A}}_{\mathbb{C}}^{N}\setminus\{0\}$.
Assume $X^{(N)}:={\mathrm{Proj\ }}{\mathcal{R}}(I)$ is normal and CM such that the restriction of $\varphi^{(N)}:X^{(N)}\to{\mathbb{A}}_{\mathbb{C}}^{N}$ onto ${\mathbb{A}}_{\mathbb{C}}^{N}\setminus\{0\}$ is either isomorphic (in case (a)) or the blowup at the smooth curve (in case (b)). Then, replacing $I$ by $I^{n}$ for $n\gg 0$,
  1. (i)
    $Z^{(N+1)}:={\mathrm{Proj\ }}{\mathcal{R}}((I,z))$ is normal and has at most one non-CM point,
  2. (ii)
    There is a projective birational morphism $g^{(N+1)}:X^{(N+1)}\to Z^{(N+1)}$ such that $X^{(N+1)}$ is normal and CM, and $Rg^{(N+1)}_{*}{\mathcal{O}}_{X^{(N+1)}}={\mathcal{O}}_{Z^{(N+1)}}$ holds.
Proof.
Let $I$ be generated by $g_{1},\ldots,g_{r}\in S_{N}={\mathbb{C}}[x_{1},\ldots,x_{N}]$, then $(I,z)\subset S_{N+1}={\mathbb{C}}[x_{1},\ldots,x_{N},z]$ is generated by $z,g_{1},\ldots,g_{r}$. By using these generators, we have an open affine covering $X^{(N)}=\bigcup_{i=1}^{r}U_{i}$, where
$U_{i}={\mathrm{Spec\ }}S_{N}\left[\frac{g_{1}}{g_{i}},\ldots,\stackrel{{\scriptstyle i}}{{\vee}},\ldots\frac{g_{r}}{g_{i}}\right].$
In the same way, we have an open affine covering $Z^{(N+1)}=\bigcup_{i=0}^{r}\widetilde{U}_{i}$, where
$\widetilde{U}_{0}={\mathrm{Spec\ }}S_{N}\left[z,\frac{g_{1}}{z},\ldots\ldots\frac{g_{r}}{z}\right]\ \ \mbox{and}$
$\begin{array}[]{lll}\widetilde{U}_{i}&={\mathrm{Spec\ }}S_{N}\left[z,\frac{z}{g_{i}},\frac{g_{1}}{g_{i}},\ldots,\stackrel{{\scriptstyle i}}{{\vee}},\ldots\frac{g_{r}}{g_{i}}\right]&\\ &={\mathrm{Spec\ }}S_{N}\left[\frac{z}{g_{i}},\frac{g_{1}}{g_{i}},\ldots,\stackrel{{\scriptstyle i}}{{\vee}},\ldots\frac{g_{r}}{g_{i}}\right]&=U_{i}\times{\mathbb{A}}_{\mathbb{C}}^{1}\ \ (i=1,\ldots,r)\\ \end{array}$
Then, for $i=1,\ldots,r$, $\widetilde{U}_{i}$ is CM since $U_{i}$ is CM. Hence, the non-CM locus on $Z^{(N+1)}$ is on the closed subscheme
$\widetilde{U}_{0}\setminus\bigcup_{i=1}^{r}\widetilde{U}_{i}=Z\left(\frac{g_{1}}{z},\ldots\ldots,\frac{g_{r}}{z}\right)=:\mathcal{Z},$
which is isomorphic to
${\mathrm{Spec\ }}{\mathbb{C}}[x_{1},\ldots,x_{N},z]/(g_{1},\ldots,g_{r})=({\mathrm{Spec\ }}{\mathbb{C}}[x_{1},\ldots,x_{N}]/I)\times{\mathbb{A}}_{\mathbb{C}}^{1}.$
Here, denote $\mathcal{Z}_{0}:={\mathrm{Spec\ }}{\mathbb{C}}[x_{1},\ldots,x_{N}]/I$, then $\mathcal{Z}_{0}$ is mapped isomorphically to the center of the blow up $\varphi^{(N)}:X^{(N)}\to{\mathbb{A}}_{\mathbb{C}}^{N}$. Let
$P:=(0,0)\in\mathcal{Z}_{0}\times{\mathbb{A}}_{\mathbb{C}}^{1}\subset\widetilde{U}_{0},$
where the first coordinate $0\in\mathcal{Z}_{0}$ corresponds to the origin $0\in{\mathbb{A}}_{\mathbb{C}}^{N}$ by $\varphi^{(N)}$, and the second coordinate $0\in{\mathbb{A}}_{\mathbb{C}}^{1}$ is the point $z=0$ in ${\mathbb{A}}_{\mathbb{C}}^{1}={\mathrm{Spec\ }}{\mathbb{C}}[z]$. Then, $P\in\widetilde{U}_{0}$ is the unique possible non-CM point in $Z^{(N+1)}$. Indeed, every point
$Q\in\mathcal{Z}_{0}\times({\mathbb{A}}_{\mathbb{C}}^{1}\setminus\{0\})$
is a smooth point in $\widetilde{U}_{0}$, because $z(Q)\neq 0$, which implies $Q$ is not in the exceptional divisor that is defined by $z=0$ on $\widetilde{U}_{0}$.
In case (a), the support of $\mathcal{Z}_{0}$ is one point. Therefore, in this case, smoothness of $\widetilde{U}_{0}$ at any point in $\mathcal{Z}_{0}\times({\mathbb{A}}_{\mathbb{C}}^{1}\setminus\{0\})$ implies that $P\in Z^{(N+1)}$ is the unique possible non-CM point.
In case (b), $\widetilde{U}_{0}$ is smooth at every point $Q\in(\mathcal{Z}_{0}\times\{0\})\setminus\{P\}$, because $Z^{(N+1)}$ is smooth on $(\psi^{(N+1)})^{-1}\left(({\mathbb{A}}_{\mathbb{C}}^{N}\setminus\{0\})\times{\mathbb{A}}_{\mathbb{C}}^{1}\right)$ by the assumption on $I$ and $Q$ is in this inverseimage. This completes the proof of (i).
Next we will find a center of the blowup to get a normal Macaulayfication as in [2, Corollary 6.12].
In case (a), let $f:=z$ and
${\mathfrak{q}}:=(z,\frac{g_{1}}{z},\ldots,\frac{g_{r}}{z})\subset{\mathfrak{m}}_{P}\subset{\mathcal{O}}_{\widetilde{U}_{0},P}.$
Then $f\in{\mathcal{O}}_{\widetilde{U}_{0},P}$ and the maximal primary ideal ${\mathfrak{q}}$ satisfy the condition of [2, Lemma 6.7], and therefore they produce a generic complete intersection $L=(y_{1},\ldots,y_{N})$, where $N=\dim Z^{(N+1)}-1$ and the locus of $L$ contains non-CM point $P\in\widetilde{U}_{0}$. (The ideal generated by the sequence obtained in [2, Lemma 6.7] is called a generic complete intersection and the generators become a generic $pS^{+}$ sequence, see [2, Remark 6.9].) By [2, Corollary 6.12], the blow-up by $\surd\overline{L}^{h}$ gives a normal and arithmetically Cohen-Macaulay for $h\gg 0$.
In case (b), there is a coordinate function, say $x_{1}$, of ${\mathbb{C}}[x_{1},\ldots,x_{N}]$ such that $x_{1}\not\in\surd\overline{I}$. Indeed, otherwise $I\subset S_{N}$ would be maximal primary at 00. Let $f:=z$ and
${\mathfrak{q}}:=(z,x_{1},\frac{g_{1}}{z},\ldots,\frac{g_{r}}{z})\subset{\mathfrak{m}}_{P}\subset{\mathcal{O}}_{\widetilde{U}_{0},P}.$
Then $f\in{\mathcal{O}}_{\widetilde{U}_{0},P}$ and the maximal primary ideal ${\mathfrak{q}}$ satisfy the condition of [2, Lemma 6.7], and therefore we also obtain a generic complete intersection $L=(y_{1},\ldots,y_{N})$. In the same way as in (a), the blow-up by $\surd\overline{L}^{h}$ gives a normal and arithmetically Cohen-Macaulay for $h\gg 0$.
Now in both cases (a) and (b), take the blowup $g^{(N+1)}:X^{(N+1)}\to Z^{(N+1)}$ by the ideal $\surd\overline{L}^{h}$, then $X^{(N+1)}$ is normal and arithmetically CM. Here, by [13, Theorem 4.1], it follows $R\varphi^{(N+1)}_{*}{\mathcal{O}}_{X^{(N+1)}}={\mathcal{O}}_{Z^{(N+1})}$. ∎
By Lemma 3.3 and 3.4, $H^{2}(X^{(N+1)},{\mathcal{O}}_{X^{(N+1)}})\neq 0$ holds if $H^{2}(X^{(N)},{\mathcal{O}}_{X^{(N)}})\neq 0$. We know $H^{2}(X^{(3)},{\mathcal{O}}_{X^{(3)}})\neq 0$. Then, by induction on the dimension $N$, we obtain for every $N\geq 3$
$\varphi^{(N)}:X^{(N)}\to{\mathbb{A}}_{\mathbb{C}}^{N}$
such that $X^{(N)}$ is normal and CM, and satisfies
$R^{i}\varphi_{*}^{(N)}{\mathcal{O}}_{X^{(N)}}|_{{\mathbb{A}}_{\mathbb{C}}^{N}\setminus\{0\}}=0,\ \ (i>0),\ \mbox{and}\ H^{2}(X^{(N)},{\mathcal{O}}_{X^{(N)}})\neq 0.$
Remark 3.5.
Let $k$ a field of charactersitic not $3$ and put $R=k[x,y,z]_{(x,y,z)}$ and ${\mathfrak{m}}=(x,y,z)$. Then Huckaba and Huneke [8, Theorem 3.11] gave an ${\mathfrak{m}}$-primary normal ideal $I\subset R$ with $H^{2}(X,{\mathcal{O}}_{X})\neq 0$, where $X={\mathrm{Proj\ }}{\mathcal{R}}(I)$. $X$ is known to be non-CM, however the dimension of non-CM locus is not known. If $k$ is an algebraically closed field and the non-CM locus of $X$ consists of finite closed points, then by starting from this, one can obtain a similar example as Example 1.5.